OpenStudy (anonymous):

Please help Verify the identity. cot^2x/csc x +1 = 1=sinx /sinx

OpenStudy (anonymous):

\[ \frac{ \cot^2x}{ \csc x +1} = \frac{ 1-sinx }{sinx} \]

OpenStudy (anonymous):

please HELP! :))))

OpenStudy (anonymous):

(Cot^2x)/(Cscx-1) = (1 - Sinx)/(Sinx) L.H.S. = (Cot^2x)/(Cscx-1) = (cosx/sinx)^2/(1/sinx - 1) L.H.S. = cos^2x/sinx(1-sinx) = (1-sin^2x)/sinx(1-sinx) L.H.S. = (1+sinx)(1-sinx)/ sinx(1-sinx) L.H.S. = (1 - sinx) / sinx = R.H.S. Is this right???

OpenStudy (pratyush5):